OpenStudy (anonymous):

For which value of a are the graphs of –4 = 3x + 6y and ax – 8y = 12 parallel? A a = –4 B a = -1/2 C a = 1/4 D a = 2

OpenStudy (freckles):

you want the slopes of both lines to be the same

OpenStudy (freckles):

so find the slopes of both lines and set them equal and solve for a

OpenStudy (anonymous):

i got -4? is that correct

OpenStudy (freckles):

I think so let's find out: \[-4=3x+6y \\ \implies -3x-4=6y \\ y=\frac{-3x}{6}-\frac{4}{6} \\ y=\frac{-1}{2}x-\frac{2}{3} \] \[ax-8y=12 \\ -8y=-ax+12 \\ y=\frac{-a}{-8}x+\frac{12}{-8} \\ y=\frac{a}{8}x+\frac{-3}{2}\] this implies that -1/2 =a/8 since the lines are parallel so yep a=-4

OpenStudy (anonymous):

i had it like that but my 4 came out as a +4

OpenStudy (freckles):

well should be -4 since -1/2 is the same as -4/8 -1/2 is not the same as 4/8

OpenStudy (freckles):

\[\frac{-1}{2}=\frac{a}{8} \\ \text{ cross multiply } \\ -1(8)=a(2) \\ -8=2a \\ \text{ divide both sides by } 2 \\ \frac{-8}{2}=\frac{2a}{2} \\ \frac{-8}{2}=a \\ a=\frac{-8}{2} \\ a=-4\]

OpenStudy (anonymous):

alot of work

OpenStudy (freckles):

you don't have to do all that work if you know that -1/2=-4/8

OpenStudy (anonymous):

oh ok i use this in other problems still

OpenStudy (anonymous):

i'll use

OpenStudy (freckles):

there is another way you could have looked at it...

OpenStudy (freckles):

\[\frac{-1}{2}=\frac{a}{8}\] you could have said what number times 2 gives me 8

OpenStudy (freckles):

and you would well miss that would 4 since 2*4=8 since you multiply the bottom number by 4 you multiply the top number by 4

OpenStudy (freckles):

\[\frac{-1}{2}=\frac{a}{8} \\ \frac{-1}{2} \cdot \frac{4}{4}=\frac{-4}{8}=\frac{a}{8} \implies a=-4\]

OpenStudy (freckles):

there is another way to think about this one and this is an easy fraction like 1 is half of 2 and what number is half of 8?

OpenStudy (freckles):

1 is half of 2 4 is half of 8 and just don't forget to involve that negative back in

OpenStudy (freckles):

but anyways do you have any questions?

OpenStudy (anonymous):

umm could u help me with this problem

OpenStudy (anonymous):

What is the slope-intercept form of the equation of the line that passes through the point (–6, 1) and is perpendicular to the graph of 2x + 3y = –5?

OpenStudy (freckles):

ok let's first determine the slope of this line

OpenStudy (freckles):

to do this we need to find the slope of 2x+3y=-5

OpenStudy (freckles):

We need to do this because we know that perpendicular lines have opposite reciprocal slopes

OpenStudy (freckles):

so if we find the slope of 2x+3y=-5 and say it is a then the slope of the line we are trying to find will be -1/a

OpenStudy (freckles):

anyways do you know how to write 2x+3y=-5 in slope-intercept form?

OpenStudy (anonymous):

i'll try

OpenStudy (anonymous):

i got 3y+2x=-5

OpenStudy (freckles):

slope-intercept form is y=mx+b

OpenStudy (freckles):

So looking at 3y+2x=-5 we are not there yet

OpenStudy (freckles):

we need to solve the equation for y

OpenStudy (freckles):

First step: Try subtracting 2x on both sides.

OpenStudy (anonymous):

ok

OpenStudy (freckles):

Let me see what you have after that step or if you were able to go further and solve for y by yourself let me see that equation pleaese

OpenStudy (anonymous):

ok im still working on the bottom part

OpenStudy (anonymous):

i keep getting 7/3 i dont know how

OpenStudy (freckles):

So you know I'm asking you to solve 3y+2x=-5 for y right? First step was to subtract 2x on both sides: This gives you 3y=-2x-5 You need to do one more step to isolate y. Do you know what it is?

OpenStudy (anonymous):

subtract the 3

OpenStudy (freckles):

no

OpenStudy (freckles):

to undo multiplication you divide

OpenStudy (freckles):

3y is 3 times y

OpenStudy (freckles):

you need to divide both sides by 3

OpenStudy (anonymous):

oh

OpenStudy (freckles):

\[3y=-2x-5 \\ \frac{3y}{3}=\frac{-2x-5}{3} \\ \\ y=\frac{-2x}{3}-\frac{5}{3} \\ y=\frac{-2}{3}x-\frac{5}{3}\] so what is the slope of this line?

OpenStudy (freckles):

The slope of y=mx+b is m The slope of y=-2/3*x-5/3 is?

OpenStudy (anonymous):

umm